Problem 21-12 Black–Scholes model

Use the Black–Scholes formula to value the following options:

a. A call option written on a stock selling for $68 per share with a $68 exercise price. The stock's standard deviation is 6% per month. The option matures in three months. The risk-free interest rate is 1.75% per month. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

b. A put option written on the same stock at the same time, with the same exercise price and expiration date. (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

(A)

S = Current Stock Price = 68
t = time until option expiration(years) = 3/12 = 0.25
X = Option Strike Price = 68
r = risk free rate(annual) = 1.75*12/100 = 0.21
s = standard deviation(annual) = 6*12/100 = 0.72
N = cumulative standard normal distribution
d1 = {ln (S/K) + (r +s^2/2)t}/s√t
= {ln (68/68) + (0.21 + 0.72^2/2)*0.25}/0.72*√0.25
= 0.3258
d2 = d1 - s√t
= 0.3258 - 0.72√0.25
= -0.0342
Using z tables,
N(d1) = 0.6277
N(d2) = 0.4864
C = Call Premium = =SN(d1) - N(d2)Ke^(-rt)
= 68*0.6277 - 0.4864*68e^(-0.21*0.25)
= 11.30

Hence, Call option premium = $11.30

(B)

S = Current Stock Price = 68
t = time until option expiration(years) = 3/12 = 0.2500
X = Option Strike Price = 68
r = risk free rate(annual) = 1.75*12/100 = 0.21
s = standard deviation(annual) = 6*12/100 = 0.72
N = cumulative standard normal distribution
d1 = {ln (S/K) + (r +s^2/2)t}/s√t
= {ln (68/68) + (0.21 + 0.72^2/2)*0.25}/0.72*√0.25
= 0.3258
d2 = d1 - s√t
= 0.3258 - 0.72√0.25
-0.0342
Using z tables,
N(-d1) = 0.3723
N(-d2) = 0.5136
P = Put Premium = =N(-d2)Ke^(-rt) - SN(-d1)
= 0.5136*68e^(-0.21*0.25) - 68*0.3723
=$7.82

Hence, Put option premium = $7.82